OpenStudy (shaleiah):

A human’s eyeball is shaped like a sphere with a diameter of 2.5 cm. A dog’s eyeball is shaped like a sphere with a diameter of 1.75 cm. About how many times greater is the volume of a human’s eyeball than the volume of a dog’s eyeball?

OpenStudy (shaleiah):

@563blackghost

563blackghost (563blackghost):

To find the volume of a sphere we would use this equation... \(\huge\color{green}{V= \frac {4}{3} \pi r^{2}}\) To find the radius we would divide the diameter by 2... \(\LARGE{2.5 \div 2=r}\) and \(\LARGE{1.75 \div 2=r}\) So what would be the radius of the human eyeball and the dog's eyeball?

OpenStudy (shaleiah):

1.25 and 0.875

563blackghost (563blackghost):

Correct ^^ Now we can find the volume of both...So we would input the radius and substitute \(\LARGE{\pi}\) as \(\Large{3.14}\).... \(\huge\color{purple}{V= \frac {4}{3}(3.14)1.25^{2}}\) this is the equation for the human eyeball.... \(\huge\color{orange}{V= \frac{4}{3}(3.14)0.875^{2}}\) this is the equation for the dog's eyeball... What would the volumes be?

OpenStudy (shaleiah):

Human=6.5416 Dog's= 3.20541

563blackghost (563blackghost):

Correct but we are gonna take it up to the hundredths place.... So we would now divide the humans eyeball by the dog's eyeball.... \(\Huge\color{red}{{6.54}\div{3.20}}\)

OpenStudy (shaleiah):

2.04

563blackghost (563blackghost):

Correct ^^ So the humans eyeball is \(\LARGE{2.04}\) times bigger than the dog's eyeball...

OpenStudy (shaleiah):

Thanks for the help!!

563blackghost (563blackghost):

\(\Huge\cal\rlap{\color{red}{No~Problem!:)}}{\color{blue}{\;No~Problem!:)}}\)